OpenStudy (anonymous):

simplify the given expression and assume that no variable equals 0. (19x^-6 y^11)(-6xy^5)

OpenStudy (kc_kennylau):

Just multiply them all while using the indices law :)

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

Huh?

OpenStudy (kc_kennylau):

\[\Large\hspace{20pt}(19x^{-6}y^{11})(-6xy^5)\]\[\Large=-6\cdot19\cdot x^{-6}\cdot x\cdot y^{11}\cdot y^5\]\[\Large=-114x^{-6+1}y^{11+5}\]\[\Large=-114x^5y^{16}\]

OpenStudy (anonymous):

how did u do that

OpenStudy (kc_kennylau):

ok i should have added one more step \[\hspace{20pt}\Large(19x^{-6}y^{11})(-6xy^5)\]\[\Large=19\cdot x^{-6}\cdot y^{11}\cdot(-6)\cdot x\cdot y^5\]\[\Large=-6\cdot19\cdot x^{-6}\cdot x\cdot y^{11}\cdot y^5\] Anyway which step was it that you didn't understand? :)

OpenStudy (anonymous):

i dont understand none of it

OpenStudy (anonymous):

how do u do this

OpenStudy (kc_kennylau):

Do you have problem with the first step

OpenStudy (anonymous):

yea

OpenStudy (kc_kennylau):

I'm just taking out all the factors

OpenStudy (kc_kennylau):

and you know the tiny little dot means multiplies?

OpenStudy (kc_kennylau):

and 19x^-6 means 19 multiplied by x^-6

OpenStudy (kc_kennylau):

So what is it that you don't understand?

OpenStudy (anonymous):

i got it now i reakon

OpenStudy (kc_kennylau):

all of it?

OpenStudy (anonymous):

no

OpenStudy (kc_kennylau):

So the second step is just regrouping

OpenStudy (anonymous):

yea

OpenStudy (kc_kennylau):

And the third step makes use of the laws of indices: \[\Large x^m\cdot x^n=x^{m+n}\]

OpenStudy (anonymous):

wat bout (12x^8 y^9)^3 divided by 8x^9 y ^18)^3

OpenStudy (kc_kennylau):

\[\Large\frac{(12x^8y^9)^3}{(8x^9y^{18})^3}\]\[\Large=\left(\frac{12x^8y^9}{8x^9y^{18}}\right)^3\]\[\Large=\left(\frac{12}8\cdot\frac{x^8}{x^9}\cdot\frac{y^9}{y^{18}}\right)^3\]\[\Large=(1.5x^{8-9}y^{9-18})^3\]\[\Large=(1.5x^{-1}y^{-9})^3\]\[\Large=3.375x^{-3}y^{-9}\]\[\Large=\frac{3.375}{x^3y^9}\] Sorry I have to sleep now please ask others or ask your teachers bye and have a good day